The Lowly Engineer
  • Metasys

    Energy Center Shift Duties 1900hrs-0700hrs Cori

    Plant rounds and readings in the Energy Center. Warmed  #3  Boiler.  Blew down  #1 Boiler and checked LWCO.  Tested boiler & Cooling tower  water and logged.   Med gas Manifold readings.  Metasys Checks and Cold and/or warm calls.

    Worked on Health stream training.

    Worked on P.M.’s

    OSC lighting turned off on cleamers they are there till midnght, turned on lights but schedule needs to be changed.

  • Metasys

    OSC Occupancy

    Released all occupancy overrides. Marked all Metasys controlled breakers & put in auto. Enabled lighting schedules & setup same as the old ASF ( M-F 6AM-7PM).

  • Metasys

    Nursing administration 1308E

    Complaints of too much air flow in nursing administration 1308E.  Area is served off of VMA 27-10.  Cfm setpoint was at over 3000.  Metastat is in room 1308-D, which has lots of computers in it.  Put override on vma 27-10 at 1500, also found d/a temp was 70 degrees on all boxes.  Checked ahu, chilled water valve not opening.  Put cooling valve controller in manual, closed butterfly halway to control d/a temp a bit warmer.   AHU is offline on Metasys due to construction.

     

    Update:  On Saturday Fred took care of Metasys issue on ahu 27-10( see post below).  Cooling valve now back in auto, override taken off of VMA 27-10.

  • Metasys

    AC27-10

    AC27-10 will be running stand alone (not on the network). We may reconnect the N2 buss next week…or not. I don’t expect any problems.

    Sorry,… the cooling valve got locked out due to an unreliable input. I added a temperature sensor to the outside air input, adjusted DAT reset & cooling lockout setpoints.

  • Metasys

    E-5012 AC 23-9 Supply Air VSD Replacement

    A replacement VSD (Danfoss Graham $ 2,772-) has been ordered for AC 23-9’s supply air.  Yesterday after the Gen Test the existing Square-D VSD did not survive.  Currently the fan is running at 100% capacity (existing VSD is completely bypassed) .  One of the end doors is intentionally propped open to relieve some of the system pressure.  Refer to req. # 7769911 for the replacement VSD status.